Your credit card charges 2% interest per month on unpaid balances. You maintain a balance of $656. How long will it take to pay off the credit card balance, assuming you make the minimum payment of $45 each month?
Answer #1

Ans 17.41 months

P = Regular Payments
PV = Loan Amount
r = rate of interest
n = no of periods
P = r (PV)
1 - (1 + r )-n
45 = 2% * 656
1 - (1 / (1 + 2%)^n))
45 = 13.12
1 - (1 / (1 + 2%)^n))
1 - (1 / (1 + 2%)^n)) = 13.12 / 45
1 - (1 / (1 + 2%)^n)) = 0.291556
n = 17.41 months